OpenStudy (anonymous):

Can someone help me solve y = 3x + 1 4y = 12x + 4 for system of equations by substitution?

OpenStudy (chaise):

y=3x+1 Eqn(1) 4y=12x+4 Eqn(2) Substitue the equation one, in to the equation 2, expand brackets and combine like terms, solving for x and then for y. 4(3x+1)=12x+4 Expanding gives: 12x+4=12x+4 It can be noted at this point, if both the sides are equal to eachother, then x=0 x=0 Substituting this value of x in to equation 1 y=3x+1 y=1